The sum of three consecutive natural numbers is 336 ineed to understand this process please break it down for me
Answer by rwm(914)

x is the first
x+1 is the second
x+2 is the third
add them
3x+3=336
3x=333
x=111
x+1=112
x+2=113
check
add them
111+112+113=336
ok